what is the slope of the tangent line of x^2 +3x at x=4 and x^2 + 3x at X=1
Answer by stanbon(75887)

what is the slope of the tangent line of x^2 +3x at x=4
f'(x) = 2x+3
f'(4) = 2*4+3 = 11
----------------------------
and x^2 + 3x at X=1
f'(x) = 3
f'(1) = 3
===============================
Cheers,
Stan H.